OpenStudy (tylermckinney16):

I got a pic posted in the comments ill fan and medals!!!!! Find the six trigonometric ratios for x. Show your work.

OpenStudy (tylermckinney16):

This is the attachment of the question.

OpenStudy (isaidavila):

Sinx=4/5 Cosx=3/5 Tgx=4/3 Ctgx=3/4 Secx=5/3 Cscx=5/4

OpenStudy (isaidavila):

Hope it helps!!

OpenStudy (tylermckinney16):

Can you explain what that is lol?

OpenStudy (isaidavila):

This are trigonometric function Sin is equal to opposite/hipothenuse Cos adyacent/hipo Tg opossite/adyacentes And the rest are the opposite Cosec inverse of sin and so on

OpenStudy (anonymous):

p = perpendicular b = base h = hypotenuse Now apply the formular for sin, cos, tan, cot, sec, cosec,. Sin = p / h Cosec = 1 / sin cos = b / h sec = 1 / cos tan = p /b cot = 1 / tan Now you know this try to put the value where it fits and find the answer.

Directrix (directrix):